# A Survey on Replication Studies and More Directions for Reasoning Language Models

### v1

@@ 3.3 Analysis and Discussions
Seed-Thinking-v1.5 (Seed, 2025) 86.7* 74.0*
Qwen3-235B (Seed, 2025) 85.7* 81.5*
@@ 3.3.1 Recipes of Training Data
Quantity and diversity are emphasized as key aspects in training reasoning language models suitable for multiple tasks. Skywork-OR1 (He et al., 2025a), Seed-Thinking-v1.5 (Seed, 2025), MiMo (Xiaomi LLM-Core Team, 2025) and Qwen3 series (Team, 2025b) have all proposed that they leverage massive RL data from various domains during training. Open-Reasoner-Zero (Hu et al., 2025) leverages data synthesis and self-distillation to expand the training dataset.
@@ 3.3.5 Reward Modeling
One of the key factors contributing to the success of RL in LLMs is the use of straightforward accuracy rewards. The minimal reward function design reduces the risk of reward hacking by leaving little room for unintended optimization. However, rule-based reward system could fail in corner cases. To address this limitation, Seed-Thinking-v1.5 (Seed, 2025) proposes a more generalizable approach that leverages LLMs to evaluate a wide range of scenarios. Their framework includes a Seed-Verifier for straightforward answer verification and a Seed-Thinking-Verifier designed for cases requiring in-depth analytical reasoning. MiMo (Xiaomi LLM-Core Team, 2025) proposes Test Difficulty Driven Reward as a sample-difficulty-aware mechanism for better accuracy rewarding.
@@ 4.1 Alternative Approaches for Reasoning Enhancement
While reinforcement learning from verifiable results (RLVR) has driven notable progress in reasoning language models, its current form remains limited, particularly in capturing intermediate reasoning steps and aligning with human expectations. To address these gaps, recent research has explored alternative approaches that complement or extend traditional RLVR techniques. In this section, we discuss two emerging directions: (i) more expressive and step-aware reward modeling methods, and (ii) preference optimization strategies that reduce computational overhead while improving training stability.
@@
Several works adopts DPO to improve the reasoning performance of language models. EXAONE Deep (Research et al., 2025) presents a series of reasoning language models which utilize 1.6M data samples for supervised fine-tuning (SFT), 20K instances for preference optimization with the method from (Xiao et al., 2025), and 10K for online RL training. As a result, EXAONE Deep 2.4B and 7.8B outperform models in comparable sizes, and EXAONE Deep 32B shows competitive performance against leading models. However, the training data is not publicly open-sourced, and there is no detail on how performance can be improved after preference optimization. Light-R1 (Wen et al., 2025a) employs a method of curriculum post-training with three training stages–two SFT stages and a DPO stage at the last. The first SFT stage uses 76k training samples, and the second one uses 3k highly difficult samples. To construct the preference pairs for the last DPO stage, it uses the rollouts from DeepSeek-R1 with verified correct answers as chosen samples, and the rollouts from the checkpoint after the second stage with verified incorrect answers as the corresponding reject samples. Iterative DPO (Tu et al., 2025) finds that DPO can rapidly improve the model’s reasoning capabilities by using various methods of constructing chosen and rejected pairs. Through multiple rounds of DPO training, they show that this method can rival the performance of online RL approaches, such as Simple-RL-Zero (Zeng et al., 2025). RedStar (Xu et al., 2025a) also studies DPO training to enhance the model’s reasoning capabilities. It constructs positive and negative samples using a rule-based reward: a reward of 1 is assigned if verification succeeds, and 0 otherwise. The study compares DPO with PPO and REINFORCE++ (Hu, 2025) and shows that DPO training is more effective than PPO in improving the model’s reasoning abilities. DPO-R1 (Zhang et al., 2025a) explores the feasibility of DPO and RAFT (Dong et al., 2023), indicating that DPO substantially enhances model performance while maintaining high training efficiency, and incorporating a SFT warm-up phase before DPO further boosts performance. Nonetheless, DPO still lags slightly behind PPO in overall effectiveness in their experiments.

### v2

@@ 3.3 Analysis and Discussions
Seed-Thinking-v1.5 (ByteDance-Seed, 2025) 86.7* 74.0*
Qwen3-235B (ByteDance-Seed, 2025) 85.7* 81.5*
@@ 3.3.1 Recipes of Training Data
Quantity and diversity are emphasized as key aspects in training reasoning language models suitable for multiple tasks. Skywork-OR1 (He et al., 2025a), Seed-Thinking-v1.5 (ByteDance-Seed, 2025), MiMo (Xiaomi LLM-Core Team, 2025) and Qwen3 series (Qwen-Team, 2025a) have all proposed that they leverage massive RL data from various domains during training. Open-Reasoner-Zero (Hu et al., 2025) leverages data synthesis and self-distillation to expand the training dataset.
@@ 3.3.5 Reward Modeling
One of the key factors contributing to the success of RL in LLMs is the use of straightforward accuracy rewards. The minimal reward function design reduces the risk of reward hacking by leaving little room for unintended optimization. However, rule-based reward system could fail in corner cases. To address this limitation, Seed-Thinking-v1.5 (ByteDance-Seed, 2025) proposes a more generalizable approach that leverages LLMs to evaluate a wide range of scenarios. Their framework includes a Seed-Verifier for straightforward answer verification and a Seed-Thinking-Verifier designed for cases requiring in-depth analytical reasoning. MiMo (Xiaomi LLM-Core Team, 2025) proposes Test Difficulty Driven Reward as a sample-difficulty-aware mechanism for better accuracy rewarding.
@@ 4.1 Alternative Approaches for Reasoning Enhancement
While reinforcement learning from verifiable rewards (RLVR) has driven notable progress in reasoning language models, its current form remains limited, particularly in capturing intermediate reasoning steps and aligning with human expectations. To address these gaps, recent research has explored alternative approaches that complement or extend traditional RLVR techniques. In this section, we discuss two emerging directions: (i) more expressive and step-aware reward modeling methods, and (ii) preference optimization strategies that reduce computational overhead while improving training stability.
@@
Several works adopt DPO to improve the reasoning performance of language models. EXAONE Deep (LG-Research et al., 2025) presents a series of reasoning language models which utilize 1.6M data samples for supervised fine-tuning (SFT), 20K instances for preference optimization with the method from (Xiao et al., 2025), and 10K for online RL training. As a result, EXAONE Deep 2.4B and 7.8B outperform models in comparable sizes, and EXAONE Deep 32B shows competitive performance against leading models. However, the training data is not publicly open-sourced, and there is no detail on how performance can be improved after preference optimization. Light-R1 (Wen et al., 2025a) employs a method of curriculum post-training with three training stages–two SFT stages and a DPO stage at the last. The first SFT stage uses 76k training samples, and the second one uses 3k highly difficult samples. To construct the preference pairs for the last DPO stage, it uses the rollouts from DeepSeek-R1 with verified correct answers as chosen samples, and the rollouts from the checkpoint after the second stage with verified incorrect answers as the corresponding reject samples. Iterative DPO (Tu et al., 2025) finds that DPO can rapidly improve the model’s reasoning capabilities by using various methods of constructing chosen and rejected pairs. Through multiple rounds of DPO training, they show that this method can rival the performance of online RL approaches, such as Simple-RL-Zero (Zeng et al., 2025). RedStar (Xu et al., 2025a) also studies DPO training to enhance the model’s reasoning capabilities. It constructs positive and negative samples using a rule-based reward: a reward of 1 is assigned if verification succeeds, and 0 otherwise. The study compares DPO with PPO and REINFORCE++ (Hu, 2025) and shows that DPO training is more effective than PPO in improving the model’s reasoning abilities. DPO-R1 (Zhang et al., 2025a) explores the feasibility of DPO and RAFT (Dong et al., 2023), indicating that DPO substantially enhances model performance while maintaining high training efficiency, and incorporating a SFT warm-up phase before DPO further boosts performance. Nonetheless, DPO still lags slightly behind PPO in overall effectiveness in their experiments.